In five years time,a father will be three times as old as his son.Four years ago the father was six times as old as his son.Find their present ages?
Answer by Maths68(1474)

Let
Present age of father = f
Present age of son = s

In five years time,
Man age = f+5
Son age = s+5
then father will be three times as old as his son.
f+5=3(s+5)
f+5=3s+15
f=3s+15-5
f=3s+10..........(1)
Four years ago
Man age = f-4
Son age = s-4
then father was six times as old as his son.
f-4=6(s-4)
f-4=6s-24
f=6s-24+4
f=6s-20..........(2)
Put the value of f from (1) to (2)
3s+10=6s-20
3s-6s=-10-20
-3s=-30
-3s/-3=-30/-3
s=10
Put the value of s in (1)
f=3(10)+10
f=30+10
f=40
Present age of father = f = 40 years old
Present age of son = s = 10 years old